# Polyhedral Complex Derivation from Piecewise Trilinear Networks

### v2

@@ B.3 Zero-trailing binarization function
Appendix C Algorithms
Here are the algorithms described in Section 4.3 and Section 5.
@@ Appendix D Theoretical proofs
$\tilde{\nu}({\bm{x}})$ $=\big{(}\nu\circ\tau\big{)}({\bm{x}})$ (33)
$=\nu\Big{(}\sum_{i=0}^{2^{D}-1}\omega(i,{\bm{w}})\cdot{\bm{H}}_{i}\Big{)}$ (34)
@@
A hypersurface $\tau({\bm{x}})=0$ intersects two points $\tau({\bm{x}}_{0})=\tau({\bm{x}}_{7})=0$ while $\tau({\bm{x}}_{1\dots 6})\neq 0$ for the remaining six points. These points form a cube, with ${\bm{x}}_{0}$ and ${\bm{x}}_{7}$ positioned on the diagonal of the cube. The hypersurface satisfies the eikonal constraint $\|\nabla\tau({\bm{x}})\|_{2}=1$ for all ${\bm{x}}\in[0,1]^{3}$. Then, the hypersurface of $\tau({\bm{x}})=0$ is a plane.
@@
By the definition of the eikonal constraint,
$\|\nabla\tau({\bm{x}})\|_{2}$ $=\Big{(}\frac{\partial\tau({\bm{x}})}{\partial x}\Big{)}^{2}+\Big{(}\frac{\partial\tau({\bm{x}})}{\partial y}\Big{)}^{2}+\Big{(}\frac{\partial\tau({\bm{x}})}{\partial z}\Big{)}^{2}=1$ (40)
$\frac{\partial\|\nabla\tau({\bm{x}})\|_{2}}{\partial x}$ $=2\Big{(}\frac{\partial\tau({\bm{x}})}{\partial y}\Big{)}\Big{(}\frac{\partial^{2}\tau({\bm{x}})}{\partial y\partial x}\Big{)}+2\Big{(}\frac{\partial\tau({\bm{x}})}{\partial z}\Big{)}\Big{(}\frac{\partial^{2}\tau({\bm{x}})}{\partial z\partial x}\Big{)}=0$ (41)
$\frac{\partial\|\nabla\tau({\bm{y}})\|_{2}}{\partial y}$ $=2\Big{(}\frac{\partial\tau({\bm{x}})}{\partial x}\Big{)}\Big{(}\frac{\partial^{2}\tau({\bm{x}})}{\partial x\partial y}\Big{)}+2\Big{(}\frac{\partial\tau({\bm{x}})}{\partial z}\Big{)}\Big{(}\frac{\partial^{2}\tau({\bm{x}})}{\partial z\partial y}\Big{)}=0$ (42)
$\frac{\partial\|\nabla\tau({\bm{z}})\|_{2}}{\partial z}$ $=2\Big{(}\frac{\partial\tau({\bm{x}})}{\partial x}\Big{)}\Big{(}\frac{\partial^{2}\tau({\bm{x}})}{\partial x\partial z}\Big{)}+2\Big{(}\frac{\partial\tau({\bm{x}})}{\partial y}\Big{)}\Big{(}\frac{\partial^{2}\tau({\bm{x}})}{\partial y\partial z}\Big{)}=0$ (43)
@@
However, since $\tau({\bm{x}}_{0})=\tau({\bm{x}}_{7})=0$ while $\tau({\bm{x}}_{2})\neq 0$ and $\tau({\bm{x}}_{5})\neq 0$, there is no solution $\tau({\bm{x}}_{1\dots 6})$ of $\frac{\partial\tau({\bm{x}})}{\partial y}=0$ for any ${\bm{x}}$. Likewise, the same goes for $\frac{\partial\tau({\bm{x}})}{\partial x}$ and $\frac{\partial\tau({\bm{x}})}{\partial z}$. By the way, we can rewrite the three partial derivatives as follows:
@@
Let $\nu_{0}({\bm{x}})$ and $\nu_{1}({\bm{x}})$ be two affine transformations defined by ${\bm{W}}_{i}^{\intercal}{\bm{x}}+{\bm{b}}_{i}$, $i\in\{0,1\}$. A hypersurface $\nu_{0}\big{(}\tau({\bm{x}})\big{)}=0$ passing two points $\nu_{0}\big{(}\tau({\bm{x}}_{0})\big{)}=\nu_{0}\big{(}\tau({\bm{x}}_{7})\big{)}=0$ while $\nu_{0}\big{(}\tau({\bm{x}}_{1\dots 6})\big{)}\neq 0$ satisfies the eikonal constraint $\|\nabla\nu_{1}\big{(}\tau({\bm{x}})\big{)}\|_{2}=1$ for all ${\bm{x}}\in[0,1]^{3}$. Then, the hypersurface is a plane.
@@ Appendix E Complexity analysis
Appendix F Experimental results
Table 2 and Table 3 show the extended outcomes from the Stanford 3D Scanning dataset in Table 1. Table 4 shows the extended outcomes from the Stanford bunny of Table 1, encompassing standard deviations, including the results considered as a pseudo-ground truth, obtained through marching cubes with 256 samples. Notice that we used the more excessive 512 samples for the pseudo-ground truth of the Medium and Large models in Figure 3 to precisely measure CD after we found that 256 samples were not enough for those cases. Table 5 shows the angular distance results. Tables 2, 3, 4 and 5 use the Small model, while Table 8 uses the Large model (ref. Section 6). We conducted all experiments with an NVIDIA V100 32GB.
@@ Appendix F Experimental results
Ours -   91147 $\pm$   203     157 $\pm$   18   6.99 53.5 $\pm$ 16.3
@@ Appendix G Visualizations
In Figure 4, we explain the rationale behind adopting the diagonal plane assumption in Theorem 4.7 and Section 4.2.

### v3

@@ B.1 Zero-trailing binarization function
B.2 Sorting polygon vertices and HashGrid marks
Algorithm 3 describes the algorithm in Section 4.3, which are sorting polygon vertices implicitly representing the direction of faces. If the viewing direction aligns opposite to the normal, the order of vertices follows a counter-clockwise arrangement. We determine the vertex order based on the normal derived from the Jacobian of the mean of vertices. We identify the counter-clockwise arrangement of vertices using dot product to compute relative angles $\theta\in[-\pi,\pi]$ and cross product for getting direction. Algorithm 4 describes the algorithm in Section 5, calculating the HashGrid marks from the hyperparameters of a HashGrid.
@@ B.2 Sorting polygon vertices and HashGrid marks
Appendix C Further discussions
C.1 Satisfying the eikonal equation and approximation
Theorem 4.7 handles where the eikonal equation does not make perfect hyperplanes. First, we observed the characteristics of trilinear hypersurfaces as illustrated in Figure 9 in Appendix. Among 64 edge scenarios within a trilinear region, a diagonal plane intersects with the hypersurface in nearly all cases. Moreover, we can take advantage of it since we can eliminate one variable by assuming the curved edge lies on the $x=z$ plane. In this light, Theorem 4.7 used linear algebra to rearrange two trilinear equations to get Equations 74 and 71. As we mentioned in the manuscript, the new vertices lie on at least two hypersurfaces, and the new edges exist on the same hypersurface, while the eikonal constraint minimizes any associated error. This error is tolerated by the hyperparameter $\epsilon$=1e-4 (Section 6) as specified in Definition 3.4.
C.2 Generalization to other positional encodings using trilinear interpolation
We can say that any encoding using trilinear interpolation (that) we defined can be applied in our method. Here, the definition is implicitly referred to Definition 4.1 (Trilinear interpolation), where the cubic corners are represented by ${\bm{H}}\in\mathbb{R}^{D\times F}$. Since we do not rely on a specific mapping function $f:\mathrm{x}\rightarrow({\bm{H}}_{0},{\bm{H}}_{1}\dots\bm{H}_{7})$, we may apply our extraction method to any encoding using trilinear interpolation that we defined. Specifically, InstantNGP [12] used a hashing function to map a corner point to the arbitrary entry of hash tables, allowing gracious collisions with online adaptivity (refer to their Sec. 3), while TensoRF [13] can be interpreted that it used a mapping function to the summation of the multiplications of two vectors from the vector-matrix factorization. Therefore, if we have the full ranks for $R_{1},R_{2},R_{3}$, in Eqn. 3 in Chen et al. [13], among the VM factorization, which can be equivalently represented by the HashGrid with a sufficient number of hash table sizes.
@@ Appendix D Theoretical proofs
$\tilde{\nu}({\bm{x}})=\big{(}\nu\circ\tau\big{)}({\bm{x}})$ $=\nu\Big{(}\sum_{i=0}^{2^{D}-1}\omega(i,{\bm{w}})\cdot{\bm{H}}_{i}\Big{)}$ (31)
@@
A hypersurface $\tau({\bm{x}})=0$ intersects two points $\tau({\bm{x}}_{0})=\tau({\bm{x}}_{7})=0$ while $\tau({\bm{x}}_{1\dots 6})\neq 0$ for the remaining six points. These points form a cube, with ${\bm{x}}_{0}$ and ${\bm{x}}_{7}$ positioned on the diagonal of the cube. The hypersurface satisfies the eikonal constraint $\|\nabla\tau({\bm{x}})\|_{2}^{2}=1$ for all ${\bm{x}}\in[0,1]^{3}$. Then, the hypersurface of $\tau({\bm{x}})=0$ is a plane.
@@
By the definitions of $\tau$ in Definition 4.3 and the eikonal constraint,
$\|\nabla\tau({\bm{x}})\|_{2}^{2}$ $=\Big{(}\frac{\partial\tau({\bm{x}})}{\partial x}\Big{)}^{2}+\Big{(}\frac{\partial\tau({\bm{x}})}{\partial y}\Big{)}^{2}+\Big{(}\frac{\partial\tau({\bm{x}})}{\partial z}\Big{)}^{2}=1$ (37)
$\frac{\partial\|\nabla\tau({\bm{x}})\|_{2}^{2}}{\partial x}$ $=2\Big{(}\frac{\partial\tau({\bm{x}})}{\partial y}\Big{)}\Big{(}\frac{\partial^{2}\tau({\bm{x}})}{\partial y\partial x}\Big{)}+2\Big{(}\frac{\partial\tau({\bm{x}})}{\partial z}\Big{)}\Big{(}\frac{\partial^{2}\tau({\bm{x}})}{\partial z\partial x}\Big{)}=0$ (38)
$\frac{\partial\|\nabla\tau({\bm{y}})\|_{2}^{2}}{\partial y}$ $=2\Big{(}\frac{\partial\tau({\bm{x}})}{\partial x}\Big{)}\Big{(}\frac{\partial^{2}\tau({\bm{x}})}{\partial x\partial y}\Big{)}+2\Big{(}\frac{\partial\tau({\bm{x}})}{\partial z}\Big{)}\Big{(}\frac{\partial^{2}\tau({\bm{x}})}{\partial z\partial y}\Big{)}=0$ (39)
$\frac{\partial\|\nabla\tau({\bm{z}})\|_{2}^{2}}{\partial z}$ $=2\Big{(}\frac{\partial\tau({\bm{x}})}{\partial x}\Big{)}\Big{(}\frac{\partial^{2}\tau({\bm{x}})}{\partial x\partial z}\Big{)}+2\Big{(}\frac{\partial\tau({\bm{x}})}{\partial y}\Big{)}\Big{(}\frac{\partial^{2}\tau({\bm{x}})}{\partial y\partial z}\Big{)}=0$ (40)
@@
However, since $\tau({\bm{x}}_{0})=\tau({\bm{x}}_{7})=0$ while $\tau({\bm{x}}_{2})\neq 0$ and $\tau({\bm{x}}_{5})\neq 0$, there is no solution $\tau({\bm{x}}_{1\dots 6})$ of $\frac{\partial\tau({\bm{x}})}{\partial y}=0$ for all ${\bm{x}}$. Likewise, the same goes for $\frac{\partial\tau({\bm{x}})}{\partial x}$ and $\frac{\partial\tau({\bm{x}})}{\partial z}$. Additionally, the three partial derivatives can be rewritten as follows:
@@
Let $\nu_{0}({\bm{x}})$ and $\nu_{1}({\bm{x}})$ be two affine transformations defined by ${\bm{W}}_{i}^{\intercal}{\bm{x}}+{\bm{b}}_{i}$, $i\in\{0,1\}$. A hypersurface $\nu_{0}\big{(}\tau({\bm{x}})\big{)}=0$ passing two points $\nu_{0}\big{(}\tau({\bm{x}}_{0})\big{)}=\nu_{0}\big{(}\tau({\bm{x}}_{7})\big{)}=0$ while $\nu_{0}\big{(}\tau({\bm{x}}_{1\dots 6})\big{)}\neq 0$ satisfies the eikonal constraint $\|\nabla\nu_{1}\big{(}\tau({\bm{x}})\big{)}\|_{2}^{2}=1$ for all ${\bm{x}}\in[0,1]^{3}$. Then, the hypersurface is a plane.
@@ Appendix E Complexity analysis
Appendix F Supplementary experiments
F.1 Complete results
Table 2 and Table 3 are extended from the Stanford 3D Scanning dataset in Table 1. Table 4 extended from the Stanford bunny of Table 1, including standard deviations, and pseudo-ground truth results, obtained through marching cubes with 256 samples. Notice that we used the more excessive 512 samples for the Medium and Large models in Figure 5. We found that 256 samples were not enough for those. Table 5 shows the angular distance results. Tables 2, 3, 4 and 5 use the Small model, while Table 8 uses the Large model (ref. Section 6). We conducted all experiments with an NVIDIA V100 32GB.
@@ F.1 Complete results
Ours -   91147 $\pm$   203     157 $\pm$   18   6.99 14.9 $\pm$ 1.66
F.2 Mesh simplification using QEM
In Figure 8, Quadric Error Metrics (QEM) [35] is a popular mesh simplification algorithm, it could potentially improve both MC and our method for efficiency. Overall, as shown, the mesh simplification favors our method. Notice that QEM entails a higher computational cost, at least $O(n\log n)$ where $n$ is the number of vertices. Additionally, these methods can cause shape deformation or holes, sensitive to hyperparameters. Therefore, QEM may be applied with caution for further mesh simplification, acknowledging the potential for unpredictable side effects. We used MeshLab’s Quadric Edge Collapse Decimation with the default settings to halve the number of vertices for each step.
F.3 Comparison with alternate approaches
Marching Cubes (MC) remains the de facto sampling-based method. In contrast, our method is a white-box method grounded in the theoretical understanding of trilinear interpolation and polyhedral complex derivation. Here is our rationale for this evaluation:
  1. 1.
    Analytical approaches for CPWA functions are limited by their inability to effectively manage spectral bias and their exponential computational cost, which struggles with exponentially growing linear regions (infeasible to run our benchmarks.) For example, the time complexity of Analytic Marching [2] is $\mathcal{O}((n/2)^{2(L-1)}|\mathcal{V}_{\mathcal{P}}|n^{4}L^{2})$, which grows exponentially with the number of layers $L$ and the width $n$ of the networks, where $|\mathcal{V}_{\mathcal{P}}|$ represents the number of vertices per face (Section 4.1 from Lei and Jia [2]). In contrast, our method has a linear time complexity with respect to the number of vertices, as discussed in Appendix E, allowing it to avoid visiting every linear region exponentially growing.
  2. 2.
    We also exclude optimization-based methods as they typically rely on MC or its variants initialization in their pipelines (our method can also be integrated into them), such as MeshSDF [36], Deep Marching Tetrahedra [10], and FlexiCubes [37]. Additionally, these methods result in prolonged computational costs.
We provided both rigorous quantitative Figure 3 and qualitative Figure 6 analyses on Marching Tetrahedra (MT) [9, 10] and Neural Dual Contour (NDC) [32], which serve as alternative representative approaches. MT is a popular variant of MC that utilizes six tetrahedra within each grid to identify intermediate vertices. NDC is a data-driven approach (a pretrained model) that uses Dual Contour. In short, MT efficiently requires SDF values but is less efficient regarding the number of vertices extracted. (MT is better than MC in the same resolution, producing more vertices, though.) NDC faces a generalization issue for a zero-shot setting, producing unsmooth surfaces. We used the code from Deep Marching Tetrahedra [10] for MT and the official code of NDC. And PyMCubes for our modernized MC library.
Figure 8: Chamfer distance for the Bunny with the Large model applying QEM (a mesh simplification). Both benefit from QEM while our method efficiently retains better chamfer distances.
F.4 Error of the learned SDF
In Equation 11 and Figure 5, we confirmed our hypothesis that the eikonal loss induces piecewise linearity in trilinear networks. To further support this, we also present the error analysis of the learned SDF. Although the Chamfer distance virtually evaluates the SDF values based on sampling, concerns may arise regarding noise in the pseudo-ground truth. For this reason, one might wish to verify whether the predicted SDFs converge toward near-zero.
We randomly sampled 100K points on the surface of pretrained networks (the Stanford bunny Large model) and calculated the average of squares for predicted SDF (model’s outputs) on the surface. As we expected, the predicted SDFs on the reconstructed surface converge to zero as the weight of the eikonal loss increases, and the number is close to zero $\approx$ 3e-8 as shown in Table 9. Note that, for a Large model, fine-grained trilinear regions from the dense grid marks (from the more dense multi-resolution grids) allow relatively small errors of 97 $\times$ 1e-9 without the eikonal loss, although the eikonal loss of 1e-2 further decreases to 31 $\times$ 1e-9. Note that the stronger eikonal loss of 1e-1 proved ineffective due to over-regularization.
Table 9: Error of the learned SDF. Weight stands for the weight of the eikonal loss and SE stands for the squared error of predicted SDFs.
Method Weight SE (1e-9) $\downarrow$
MC64 1e-2 2308
MC256 1e-2     66
Ours 0     97
1e-4     91
1e-3     68
1e-2     31
F.5 Effect of model sizes
In Table 10, we explore the ablation study to determine whether model size impacts model fitting, which in turn affects the GT mesh (we used MC256 for the pseudo-ground truth of the Small models). With a smaller model size, the learned mesh becomes simpler due to underfitting. As our evaluation focuses on how accurately the extraction methods capture the underlying zero level-set decision boundary of the networks, the change in CD remains small.
Table 10: Ablation study varying the number of layers and the width of networks for the Stanford bunny Small models. And, Note indicates the architecture of the decoding networks.
Layers Width Vertices CD (1e-6) $\downarrow$ Time (sec) $\downarrow$ Note
2   4 4565 738 0.11
2   8 4863 768 0.13
2 16 4566 734 0.12
2 32 4577 754 0.22
2 64 4611 739 0.50 InstantNGP [12]
3   4 4547 666 0.12
3   8 4661 750 0.19
3 16 4628 724 0.65 Ours
3 32 4567 754 0.57
3 64 5588 759 2.86
Notice that the model with the number of layers of 3 and the width of 64 has the longest runtime of 2.80, primarily because the number of intermediate vertices exceeds 80K. The time complexity is more sensitive to the number of vertices than to the number of layers or the width, which are only indirect factors. Note that the model capacity of the decoding networks is comparable to InstantNGP [12], where the representational efficiency is driven by the HashGrid.
@@ Appendix G Visualizations
For Figure 6 in the main paper (closer looks at bunny’s nose), we want to leave a note about the small facets along the left-side of the nose line in the highlighted region. We assume these narrow facets are embedded in the facets of MC256 (pseudo-ground truth), while MC64 fails to capture the left and right edge of the nose and represents this with more facets cutting these two sides of edges – not aligned with nose lines and rather assign more facets in here.
In Figure 9, we explain the rationale behind adopting the diagonal plane assumption in Theorem 4.7 and Section 4.2. We illustrated 64 edge scenarios within a trilinear region, omitting isomorphic cases.